OpenStudy (anonymous):

Evaluate X from the equation log 4x + log X=0.8 2 2

OpenStudy (anonymous):

use the law of logarithms that says \[\log _{a}x+\log _{a}y=\log _{a}xy\] to combine the left side of equation \[\log _{2}4x*x=.8\] any idea where to go from here?

OpenStudy (anonymous):

thanks for the reply, yeah I got to that step by simplifying and using the proper tie for have put above but I wasn't sure what to do.